# Observing radio transients with Phased ALMA: Pulses from the Galactic Centre magnetar

### v2

@@ Abstract
Context. Radio transients, such as pulsars and Fast Radio Bursts (FRBs), are primarily detected at centimetre (cm) radio wavelengths, where higher luminosities are found. However, observations of sources in dense environments are heavily affected by propagation effects, such as scattering, which may hinder a detection. Millimetre (mm)-wave observations bypass this complication but require the largest radio telescopes to compensate for the lower flux densities. When used in phased mode, the ALMA radio telescope provides an equivalent dish size of 84m, making it the most sensitive instrument at mm/sub-mm wavelengths. In combination with its high-time resolution, it offers a unique opportunity to study radio transients in an unexplored frequency window.
Aims. We study the Galactic Centre (GC) magnetar, PSR J1745-2900, as a laboratory for magnetars in complex magneto-turbulent environments and for linking with FRBs. Through this pilot study, we showcase the potential of ALMA in its phased configuration to observe radio transients and to achieve, for some sources, the first ever detections outside the cm-wave range.
Methods. We studied the GC magnetar using ALMA archival data of Sgr A* at Band-3, taken during the 2017 GMVA campaign. The data were searched in intensity, and the pulses were classified based on their circular and linear polarisation properties and arrival phase.
@@
A particular type of radio emitting neutron stars (NSs) are magnetars. They are young NSs ( yr ; Kouveliotou 1999) with extremely strong magnetic fields, in the order of (Duncan & Thompson, 1992). Although more than thirty magnetars have been reported (Olausen & Kaspi, 2014), only a handful exhibit detectable emission at radio frequencies. Recently, they have attracted significant attention as promising candidates for the origin of Fast Radio Bursts (FRBs; Masui et al. 2015, Michilli et al. 2018). FRBs were first discovered while inspecting archival data of pulsar surveys (Lorimer et al., 2007). Since then, over 800 FRBs have been discovered (Xu et al., 2023), with the majority classified as one-off events and only 5% as repeaters (Petroff et al., 2022). These radio transient events, of extragalactic origin, are characterized by their high inferred luminosities (, Locatelli et al. 2019) and pulses as short as microseconds (Snelders et al., 2023). Although there is no widely accepted explanation for the origin of FRBs, one of the strongest evidence of them being magnetars at cosmological distances is the detection of two bright radio bursts from the Galactic magnetar SGR 1935+2154. These bursts exceeded the typical isotropic energy of regular radio magnetars by approximately three orders of magnitude (Bochenek et al., 2020; Kirsten et al., 2021), but are still relatively weak compared to FRBs.
Commonly, the observations of radio transient sources, such as pulsars and FRBs, are performed using large radio telescopes operating at centimetric wavelengths. This is because these wavelengths yield higher flux densities (Manchester & Taylor, 1977). Pulsars, for instance, are known to have a steep spectral index around = -1.6 (Jankowski et al., 2017a). However, some magnetars have flatter or inverted values (e.g., Torne et al. 2015, 2017). The detectable flux densities are also affected by interstellar effects such as scintillation, scattering and dispersion. The latter introduces a frequency-dependent delay in the arrival of the signal, given by:
where s, and are the lower and higher frequencies of the observing band, respectively (Lorimer & Kramer, 2004). DM refers to the dispersion measure, and quantifies the integrated column density of electrons along the line of sight:
@@
From Eq. 1, by fixing the DM, it is evident that the lower the frequency, the higher the dispersion delay. In particular, for a source with = 1770 , observing at 1.4 GHz the delay is approximately 1.793 s, while at 42 GHz is 366 s, and at 86 GHz is 46 s (see Table 1).
@@
ALMA (Band-1) | 42.168 | 44.168 | 43.168 | 2000 | 32 | 366 | 11 |
ALMA (Band-3) | 85.268 | 87.268 | 86.268 | 2000 | 32 | 46 | 1 |
@@
Due to its complex and scattered medium, the vicinity of Sgr A∗ holds the potential to be used as a laboratory for testing the behaviour of pulsars and magnetars in extreme environments and to link them with FRBs. The DM excess and highly varying rotation measure (RM) of FRBs indicate that they could be born in complex magneto-turbulent environments (Platts et al., 2019; Mottez et al., 2020; Wang et al., 2022). Similar properties could be found around pulsars and the magnetar near the Galactic Centre (GC, Desvignes et al. 2018, Abbate et al. 2023). For instance, large RMs are expected in the boundaries of massive black holes (Bower et al., 2003; Michilli et al., 2018). Importantly, given the high plasma densities of regions such as the GC (; Witzel et al. 2021), propagation effects such as the dispersion delay and scattering may hinder a detection at low radio frequencies (Torne et al., 2023). For scattering, the timescale is inversely proportional to the fourth power of the frequency as (Lorimer & Kramer, 2004). One way to overcome these two propagation effects and reach the densest regions is to observe at higher frequencies. To compensate for the lower flux densities, the largest possible radio telescopes are required.
@@
The ALMA Phasing System (APS, Matthews et al. 2018) was developed for Very Long Baseline Interferometry (VLBI) observations in global campaigns such as the Global Millimeter VLBI Array (GMVA) and the Event Horizon Telescope (EHT, Akiyama et al. 2019). APS is available with an odd number of antennas as a hardware requirement (Matthews et al., 2018). First it was implemented in Band-6 and Band-3 (Goddi et al., 2019) and then at Band-7 (Crew et al., 2023) and Band-1. When combining 49 of its 12m antenna, its equivalent dish size is 84m, which makes it the most sensitive instrument at mm/sub-mm wavelengths. Previous observations of known pulsars with phased ALMA have detected sources such as the Vela Pulsar (Liu et al., 2019). Searches for new pulsars have been also conducted in the recent years (Liu et al., 2021; Torne et al., 2023). For FRBs, thus far, no observations have been performed with phased ALMA.
@@
Based on APS, the Phased ALMA mode (PAM) was introduced in Cycle 8 (2021) to enable observations of weak radio sources (¡50 mJy) in time domain. The phasing of ALMA antennas involves multiple steps (Goddi et al., 2019). First, phase corrections align each signal to a reference antenna. Then, the signals of each antenna are coherently summed and the raw voltage data are recorded as 2-bit samples using MARK 6 recorders. When the output is used for pulsar searching, the data are packetized and divided in frequency channels, with a time resolution multiple of 8 s (Cortes et al., 2024).
@@
In this work we studied the GC magnetar, PSR J1745-2900, using ALMA archival data of Sgr A∗. This is because the source is located within the same synthetized beam of Sgr A∗ ( 5.7”), due to its 0.07-2 pc projected distance (Rea et al., 2013). We searched for the magnetar’s single pulses to link them to FRBs emission. We examined the potential of phased ALMA to open a window to higher frequency observations of radio transients. This paper is organized as follows: in Section 2 we describe the observations, then in Section 3 the single pulse search. In Section 4 we present the results of the pulses characterization. In Section 5 we discuss the future prospects of using phased ALMA for radio transient observations. Finally, in Section 6 we present our final remarks.
@@ 2 Observations
We used observations of Sgr A* from the 2017 GMVA campaign at 3.5 mm. The observations were carried out on April 3, using 37 12-m antennas in Band-3 configuration centred at 86 GHz (Liu et al., 2021). The dataset had a total of 5.2 hrs, with on-source time of 2.5 hrs. We used a total of 2 hr observing time, excluding 30 minutes of unusable data (see Table 2). The off-source time corresponds to the calibrations needed between source scans. Some calibrators were 3C 279, NRAO 530 and J1924-2914 (OV-236).
As discussed in Liu et al. (2021), the data was converted to PSRFITS format (Hotan et al., 2004) with full Stokes information (I,Q,U,V) and 32 frequency channels, each with a 62.5 MHz channel width. The outputs were multiple files corresponding to each scan, with an approximate duration of 52 s. In every 18.192 s there was a drop-off on the timeseries caused by the phasing cycle of the array. This artifact was previously reported in Liu et al. (2021). We listed the observations in Table 2.
@@ 3 Methods and Data Processing
We searched for single pulses using a PRESTO (Ransom, 2011) based pipeline, detailed in Braga et al. (2025) with some modifications to account for PSRFITS containing Stokes. We used a signal-to-noise ratio (SNR) threshold of 6, determined through analysis of the noise floor of the time series on pulse-free regions. These time series were constructed by summing all the frequency channels of the PSRFITS files on Stokes I. Given the high DM of the source and the high observing frequency at Band-3, the dispersion delay between the highest and the lowest frequency channel is 46 s (see Table 1). Therefore, DM trials were unnecessary, and we adopted the reported value of the source, 1770 (Liu et al., 2021). Using this value we searched for single pulses in Stokes I. We inspected each candidate visually. To this end, we constructed the time series in intensity, linear and circular polarisation.
Importantly, due to the negligible dispersion delay, we rely primarily on polarisation to confirm the pulse candidates. We complemented with a second test, in which we use the spin period of the magnetar to assign a phase to the arrival time of each pulse. This arrival phase is compared with the region where the pulse is expected to arrive. Such region was determined from the integrated pulse profile width 333We define the width of the integrated pulse profile as the full-width until the SNR drops to the noise floor.. We used this method as most single pulses are significantly narrower than their integrated pulse, which is built up from adding up thousands of pulses (Drake & Craft, 1968). Nevertheless, in some cases magnetars also show single pulses outside the main profile (Camilo et al., 2007).
According to our criteria, if a pulse candidate had significant linear and circularly polarised flux and it arrived in phase with the integrated pulse profile of PSR J17452900, then it is considered as a real pulse. It is worth noting that pulses with small polarisation fractions would have not been detected by our pipelines.
@@ 4 Results
We report eight single pulses from the GC magnetar, PSR J1745-2900, which are shown in Fig. 1. The bottom panel shows the dynamic spectra across all 32 frequency channels. The top panel displays the integration along the bandwidth, we also refer to this as the single pulse profile. We perform this integration using Stokes I. The pulses, together with their linearly and circularly polarized flux, are shown with respect to their arrival phase in Fig. 2. We calculate the times of arrival (TOAs) of each reported pulse. In our search we found the three pulses reported by Liu et al. (2021) plus five new ones.
Using all data blocks from Table 2, we create integrated pulse profiles per scan. From the ones with highest SNR, we obtain TOAs that we input to TEMPO (Nice et al., 2015), following standard timing procedures. This step was necessary as the ephemeris listed in the ATNF catalogue (Manchester et al., 2005) was not sufficient to fully recover the signal of the magnetar. Through residual minimisation, with the spin period as a the only free parameter, we generate a new ephemeris file shown in Table 3.
@@
Right ascension, (J2000) | 14:45:40.1662(8) |
Declination, (J2000) | 29:00:29.89(1) |
Spin frequency, () | 0.26762(8) |
Dispersion measure, DM () | 1770(3) |
Reference Epoch, PEPOCH | 57846.349663 |
@@ 4.1 Pulse characterisation
In Table 4 we report the pulses’ properties. We report the modified Julian day (MJD), SNR, full width at half maximum (FWHM), flux density, fluence 555Corresponding to the band-averaged peak flux density. and isotropic energy. As it is seen in Fig. 1, most of the pulses from the GC magnetar have more than one component. We compute the SNR by normalizing the time series using a pulse-free region and taking the pulse peak. To estimate the pulse width, we use a multi-Gaussian fitting routine developed for single pulse analysis (Limaye et al. in prep.). Then, the fluence is obtained as:
where is the number of polarisations, the number of spectral windows and is the bandwidth of each spectral window. The system equivalent flux density (SEFD) is a parameter given for each telescope as:
where is the Boltzmann constant, is the system temperature, G is the system gain of phased ALMA (51 K and 1.15 K/Jy respectively on Liu et al. 2021) and the effective area is . Then, the isotropic energy is given by:
where L corresponds to the source luminosity distance, equal to 8.3 kpc (Yao et al., 2017). As it is seen in Table 4, the energy of the pulses is in the order of erg.
Candidate | SNR | MJD (bary) | FWHM (s) | Flux density (Jy) | Fluence (Jy ms) | Energy (1029 erg) |
1 | 11.0 | 57846.3495 | 0.11 | 0.023 | 2.589 | 4.27 |
2 | 6.6 | 57846.5413 | 0.13 | 0.013 | 1.660 | 2.74 |
3 | 15.6 | 57846.5730 | 0.04 | 0.058 | 2.068 | 3.41 |
4 | 22.2 | 57846.5422 | 0.04 | 0.074 | 3.284 | 5.41 |
5 | 7.5 | 57846.3478 | 0.08 | 0.018 | 1.499 | 2.47 |
6 | 10.7 | 57846.3493 | 0.04 | 0.036 | 1.556 | 2.57 |
7 | 8.9 | 57846.5412 | 0.10 | 0.020 | 1.956 | 3.23 |
8 | 14.7 | 57846.3767 | 0.01 | 0.127 | 0.837 | 1.38 |

### v4

@@ Abstract
Context. Radio transients, such as pulsars and fast radio bursts (FRBs), are primarily detected at centimetre (cm) radio wavelengths, where the highest luminosities are found. However, observations of sources in dense environments are heavily affected by propagation effects, such as scattering, which may hinder a detection. Millimetre (mm)-wave observations bypass this complication but require the largest radio telescopes to compensate for the lower flux densities. When used in phased mode, the ALMA radio telescope provides an equivalent dish size of 84m, making it the most sensitive instrument at mm/sub-mm wavelengths. In combination with its high time resolution, it offers a unique opportunity to study radio transients in an unexplored frequency window.
Aims. We studied the Galactic Centre (GC) magnetar, PSR J1745-2900, as a laboratory for magnetars in complex magneto-turbulent environments and for linking with FRBs. Through this pilot study, we showcase the potential of ALMA in its phased configuration to observe radio transients and to achieve, for some sources, the first ever detections outside the cm-wave range.
Methods. We studied the GC magnetar using ALMA archival data of Sgr A* at Band 3, taken during the 2017 GMVA campaign. The data were searched in intensity, and the pulses were classified based on their circular and linear polarisation properties and arrival phase.
@@
A particular type of radio emitting neutron stars (NSs) are magnetars. They are young NSs ( yr ; Kouveliotou 1999) with extremely strong magnetic fields, of the order of (Duncan & Thompson 1992). Although more than 30 magnetars have been reported (Olausen & Kaspi 2014), only a handful exhibit detectable emission at radio frequencies. Recently, they have attracted significant attention as promising candidates for the origin of fast radio bursts (FRBs; Masui et al. 2015, Michilli et al. 2018). FRBs were first discovered while inspecting archival data of pulsar surveys (Lorimer et al. 2007). Since then, over 4000 FRBs have been discovered (Siegel & Fonseca 2025), the majority classified as one-off events and only 5% as repeaters (Petroff et al. 2022). These radio transient events, of extragalactic origin, are characterised by their high inferred luminosities (, Locatelli et al. 2019) and pulses as short as microseconds (Snelders et al. 2023). Although there is no widely accepted explanation for the origin of FRBs, one of the strongest pieces of evidence that they could be magnetars at cosmological distances is the detection of two bright radio bursts from the Galactic magnetar SGR 1935+2154. These bursts exceeded the typical isotropic energy of regular radio magnetars by approximately three orders of magnitude (Bochenek et al. 2020; Kirsten et al. 2021), but are still relatively weak compared to FRBs.
Commonly, the observations of radio transient sources, such as pulsars and FRBs, are performed using large radio telescopes operating at centimetre (cm) wavelengths. This is because these wavelengths yield higher flux densities (Manchester & Taylor 1977). Pulsars, for instance, are known to have a steep spectral index around = -1.6 (Jankowski et al. 2017a). However, some magnetars have flatter or inverted values (e.g., Torne et al. 2015, 2017). The detectable flux densities are also affected by interstellar effects such as scintillation, scattering and dispersion. The last of these effects introduces a frequency-dependent delay in the arrival of the signal, given by
where s, and are the lower and higher frequencies of the observing band, respectively (Lorimer & Kramer 2004). DM refers to the dispersion measure, and quantifies the integrated column density of electrons along the line of sight
@@
From Eq. 1, by fixing the DM, it is evident that the lower the frequency, the higher the dispersion delay. In particular, for a source with = 1770 , observing at 1.4 GHz the delay is approximately 1.793 s, while at 42 GHz it is 366 s, and at 86 GHz it is 46 s (see Table 1).
@@
| ALMA (Band 1) | 42.168 | 44.168 | 43.168 | 2000 | 32 | 366 | 11 |
| ALMA (Band 3) | 85.268 | 87.268 | 86.268 | 2000 | 32 | 46 | 1 |
@@
Due to its complex and scattered medium, the vicinity of Sgr A∗ holds the potential to be used as a laboratory for testing the behaviour of pulsars and magnetars in extreme environments and the ability to link them with FRBs. The DM excess and highly varying rotation measure (RM) of FRBs indicate that they could be born in complex magneto-turbulent environments (Platts et al. 2019; Mottez et al. 2020; Wang et al. 2022). Similar properties could be found around pulsars and the magnetar near the Galactic Centre (GC, Desvignes et al. 2018, Abbate et al. 2023). For instance, large RMs are expected in the boundaries of massive black holes (Bower et al. 2003; Michilli et al. 2018). Importantly, given the high plasma densities of regions such as the GC (; Witzel et al. 2021), propagation effects such as the dispersion delay and scattering may hinder a detection at low radio frequencies (Torne et al. 2023). For scattering, the timescale is inversely proportional to the fourth power of the frequency as (Lorimer & Kramer 2004). One way to overcome these two propagation effects and reach the densest regions is to observe at higher frequencies. To compensate for the lower flux densities, the largest possible radio telescopes are required.
@@
The ALMA Phasing System (APS, Matthews et al. 2018) was developed for very long baseline interferometry (VLBI) observations in global campaigns such as the Global Millimetre VLBI Array (GMVA, Ros et al. 2024) and the Event Horizon Telescope (EHT, Akiyama et al. 2019). APS is available with an odd number of antennas as a hardware requirement (Matthews et al. 2018). First it was implemented in Band 6 and Band 3 (Goddi et al. 2019) and then at Band 7 (Crew et al. 2023) and Band 1. When combining 49 of its 12m antennas, its equivalent dish size is 84m, which makes it the most sensitive instrument at millimetre (mm) to sub-millimetre (sub-mm) wavelengths. Previous observations of known pulsars with phased ALMA have detected sources such as the Vela Pulsar (Liu et al. 2019). Searches for new pulsars have also been conducted in recent years (Liu et al. 2021; Torne et al. 2023). For FRBs, no observations have been performed with phased ALMA to date.
@@
Based on APS, the Phased ALMA mode (PAM) was introduced in Cycle 8 (2021) to enable observations of weak radio sources (¡50 mJy) in time domain. The phasing of ALMA antennas involves multiple steps (Goddi et al. 2019). First, phase corrections align each signal to a reference antenna. Then, the signals of each antenna are coherently summed and the raw voltage data are recorded as 2 bit samples using MARK 6 recorders. When the output is used for pulsar searching, the data are packeted and divided in frequency channels, with a time resolution multiple of 8 s (Cortes et al. 2024).
@@
For this work we studied the GC magnetar, PSR J1745-2900, using ALMA archival data of Sgr A∗. This is because the source is located within the same synthesised beam of Sgr A∗ ( 5.7”), due to its 0.07-2 pc projected distance (Rea et al. 2013). We searched for the magnetar –single pulses to link them to FRB emission. We examined the potential of phased ALMA to open a window to higher frequency observations of radio transients. This paper is organized as follows. In Section 2 we describe the observations, and then in Section 3 the single pulse search. In Section 4 we present the results of the pulses characterisation. In Section 5 we discuss the future prospects of using phased ALMA for radio transient observations. Finally, in Section 6 we present our final remarks.
@@ 2 Observations
We used observations of Sgr A* from the 2017 GMVA campaign at 3.5 mm (Project Code: 2017.1.00795.V, PI: M. Johnson). The observations were carried out on April 3, using 37 12m antennas in Band 3 configuration centred at 86 GHz (Liu et al. 2021). The dataset covered a total of 5.2 hr, with on-source time of 2.5 hr. We used a total of 2 hr observing time, excluding 30 minutes of unusable data (see Table 2). The off-source time corresponds to the calibrations needed between source scans. Some calibrators were 3C 279, NRAO 530 and J1924-2914 (OV-236).
As discussed in Liu et al. (2021), the data was converted to PSRFITS format (Hotan et al. 2004) with full Stokes information (I,Q,U,V) and 32 frequency channels, each with a 62.5 MHz channel width. The outputs were multiple files corresponding to each scan, with an approximate duration of 52 s. Every 18.192 s there was a drop-off on the time series caused by the phasing cycle of the array. This artefact was previously reported in Liu et al. (2021). We list the observations in Table 2.
@@ 3 Methods and data processing
We searched for single pulses using a PRESTO-based pipeline (Ransom 2011), detailed in Braga et al. (2025) with some modifications to account for PSRFITS containing Stokes. We used a signal-to-noise ratio (S/N) threshold of 6, determined through analysis of the noise floor of the time series on pulse-free regions. These time series were constructed by summing all the frequency channels of the PSRFITS files on Stokes I. Given the high DM of the source and the high observing frequency at Band 3, the dispersion delay between the highest and the lowest frequency channel is 46 s (see Table 1). Therefore, DM trials were unnecessary, and we adopted the reported value of the source, 1770 (Liu et al. 2021). Using this value we searched for single pulses in Stokes I. We inspected each candidate visually. To this end, we constructed the time series in intensity, linear and circular polarisation.
Importantly, due to the negligible dispersion delay, we relied primarily on polarisation to confirm the pulse candidates. We complemented these data with a second test –in which we used the spin period of the magnetar to assign a phase to the arrival time of each pulse. This arrival phase was compared with the region where the pulse is expected to arrive. This region was determined from the full width of the integrated pulse profile. We used this method as most single pulses are significantly narrower than their integrated pulse, which is built up from adding up thousands of pulses (Drake & Craft 1968). Nevertheless, in some cases magnetars also show single pulses outside the main profile (Camilo et al. 2007).
According to our criteria, if a pulse candidate has significant linear and circularly polarised flux and it arrives in phase with the integrated pulse profile of PSR J17452900, then it is considered a real pulse. It is worth noting that pulses with small polarisation fractions would have not been detected by our pipelines.
@@ 4 Results
We report eight single pulses from the GC magnetar, PSR J1745-2900, which are shown in Fig. 1. The bottom panel shows the dynamic spectra across all 32 frequency channels. The top panel displays the integration along the bandwidth, we also refer to this as the single pulse profile. We performed this integration using Stokes I. The pulses, together with their linearly and circularly polarized flux, are shown with respect to their arrival phase in Fig. 2. We calculated the time of arrival (TOA) of each reported pulse. In our search we found the three pulses reported by Liu et al. (2021) plus five new pulses.
Using all data blocks from Table 2, we created integrated pulse profiles per scan. From those with the highest S/N, we obtained TOAs that we input to TEMPO (Nice et al. 2015), following standard timing procedures. This step was necessary as the ephemeris listed in the ATNF catalogue (Manchester et al. 2005) was not sufficient to fully recover the signal of the magnetar. Through residual minimisation, with the spin period as the only free parameter, we generated a new ephemeris file shown in Table 3.
@@
| Right ascension, (J2000). | 17:45:40.1662(8) |
| Declination, (J2000). | 29:00:29.89(1) |
| Spin frequency, (). | 0.26762(8) |
| Dispersion measure, DM (). | 1770(3) |
| Reference Epoch, PEPOCH . | 57846.349663 |
@@ 4.1 Pulse characterisation
In Table 4 we report the pulse properties. We report the modified Julian day (MJD), S/N, full width at half maximum (FWHM), flux density, fluence 444Corresponding to the band-averaged peak flux density. and isotropic energy. As shown in Fig. 1, most of the pulses from the GC magnetar have more than one component. We computed the S/N by normalising the time series using a pulse-free region and taking the pulse peak. To estimate the pulse width, we used a multi-Gaussian fitting routine developed for single pulse analysis (Limaye et al. in prep.). Then, the fluence is obtained as
where is the number of polarisations, the number of spectral windows and the bandwidth of each spectral window. The system equivalent flux density (SEFD) is a parameter given for each telescope as
where is the Boltzmann constant, is the system temperature, G is the system gain of phased ALMA (51 K and 1.15 K/Jy respectively on Liu et al. 2021), and is the effective area. Then, the isotropic energy is given by
where D corresponds to the source luminosity distance, equal to 8.3 kpc (Yao et al. 2017). As listed in Table 4, the energy of the pulses is of the order of erg.
| Candidate | S/N | MJD (bary) | FWHM (s) | Flux density (Jy) | Fluence (Jy ms) | Energy (1029 erg) |
| 1 | 11.0 | 57846.3495 | 0.11 | 0.023 | 2.589 | 4.27 |
| 2 | 6.6 | 57846.5413 | 0.13 | 0.013 | 1.660 | 2.74 |
| 3 | 15.6 | 57846.5730 | 0.04 | 0.058 | 2.068 | 3.41 |
| 4 | 22.2 | 57846.5422 | 0.04 | 0.074 | 3.284 | 5.41 |
| 5 | 7.5 | 57846.3478 | 0.08 | 0.018 | 1.499 | 2.47 |
| 6 | 10.7 | 57846.3493 | 0.04 | 0.036 | 1.556 | 2.57 |
| 7 | 8.9 | 57846.5412 | 0.10 | 0.020 | 1.956 | 3.23 |
| 8 | 14.7 | 57846.3767 | 0.01 | 0.127 | 0.837 | 1.38 |